OpenStudy (anonymous):

help What is the graph of the system? y ≤ -x – 1 y ≥ 2x + 4 asap

OpenStudy (bahrom7893):

just graph: y=-x-1 and y=2x+4 and then shade

OpenStudy (bahrom7893):

They're both solid btw, since u have \[\ge\] and \[\le\]

OpenStudy (anonymous):

shade which way

sam (.sam.):